# Demystifying Misconceptions in Social Bots Research

### v2

@@ Introduction
Some work in the literature fails to follow this methodology, only to accuse other researchers of bias. Gallwitz and Kreil (2022) demonstrate this behavior in several cases. Without disclosing any annotation rubric, they change criteria for labeling bots multiple times, thus inflating the false-positive or false-negative rates of bot detection methods to fit their narrative. In one case, criticizing Yang et al. (2019), they label accounts that automatically post news headlines, such as those associated with major newswire agencies, as human. Their criticism of Keller and Klinger (2019) goes further: first, Gallwitz and Kreil considered a different set of accounts than those analyzed by Keller and Klinger; second, they manually inspected only 0.04% of this different set; third, they ignored suspended accounts; and, fourth, they labeled as human accounts those that cross-posted tweets through software apps. Their conclusion was that «not a single one of the accounts in [Keller and Klinger’s] list was a “bot” in any meaningful sense of the word.» Similarly biased annotations are used by Gallwitz and Kreil (2022) to criticize Dunn et al. (2020): suspended accounts are again assumed to be human to boost the alleged false-positive rate, then accounts that are clearly automated are still labeled as classification errors by changing the standard from “bot” to “malicious” or “spreading vaccine-critical information.” The use of obsolete or biased datasets with unclear, inconsistent, or shifting definitions is a dangerous trend toward scientific manipulation.
@@
The assumption about the possibility to build comprehensive bot datasets is equally flawed. Creating such datasets would require a uniform and random sample of an adequate number of accounts from the complete distribution of existing bots. This would allow obtaining an accurate representation of the bot landscape that captures the full complexity of the problem space. While in some domains of application for machine learning this is feasible, [^2] For example, in certain speech recognition tasks all instances (e.g., languages) to be classified are known in advance. Given a sufficiently large and diverse set of user inputs, unbiased and comprehensive datasets are possible (Ardila et al., 2020). we argue that performing a uniform and random sample of the whole population of bots is extremely problematic—if not outright impossible. In addition to the many challenges brought by social data (Olteanu et al., 2019; Pasquetto et al., 2020), the task of detecting social bots also partly belongs to the fields of information security and open-source intelligence, which are intrinsically characterized by the presence of adversaries who are strongly motivated to remain hidden (Cresci et al., 2021; Torres-Lugo et al., 2022). Indeed, there has always been limited knowledge of the real extent of the bot problem, which makes it hard to track bots on a platform, quantify their numbers, and assess their impact (Varol, 2023; Mendoza et al., 2020; Tan et al., 2023). We only have a partial understanding and visibility of the botnets that operate on our online platforms. [^3] The court dispute between Elon Musk and former Twitter CEO Parag Agrawal on the real number of bots on Twitter is a paramount example of the intrinsic difficulties and ambiguity of tracking social bots. Figure 1 provides a conceptual view of the landscape of social bots and human-operated accounts: firstly according to a naïve interpretation (Figure 1(a)) and then in a more realistic representation (Figure 1(b)). Figure 1(b) highlights the knowledge gaps related to the presence of bots in online platforms. These include known unknowns: the exact number of bots belonging to known botnets. But they also include unknown unknowns: information about the hidden botnets operating on a platform. How can we create unbiased and comprehensive datasets when we largely ignore the types and numbers of bots populating online platforms? This fundamental question cannot currently be answered.
@@
Part of the social bot literature claims that the bot detection task has been effectively solved. Proponents of this perspective highlight the perceived accuracy and reliability of current detection methods, and advocate their suitability for integration into various downstream applications. This optimistic view positions bot detection research as a success story for machine learning, highlighting the progress made in achieving practical and deployable solutions in the field.
@@
We disagree with this interpretation of the state of social bot research. On the contrary, we believe that unbiased and responsible research in this area should acknowledge the body of literature warning researchers and practitioners about the limitations of bot detection. We are well aware of this body of work, having actively contributed to its development for many years. Below we report on some notable examples. The first accounts of the difficulties supervised detectors face in detecting sophisticated bots date back to 2013 (Yang et al., 2013) and intensified in subsequent years (Cresci et al., 2017; Grimme et al., 2018; Cresci et al., 2019). The limitations of benchmark datasets for bot detection were discussed by Feng et al. (2021) and the data-related limitations to the generalizability and replicability in social bots research were touched upon by Assenmacher et al. (2022). Several works measured and discussed the generalization deficiencies of bot detectors, with particular emphasis on supervised ones (Echeverría et al., 2018; Yang et al., 2019; Cresci, 2020; Sayyadiharikandeh et al., 2020; Yang et al., 2020; Feng et al., 2021). Others warned about the lack of consistency in bot classifications obtained at different points in time (Rauchfleisch and Kaiser, 2020) and with different detectors (Svenaeus, 2020; Martini et al., 2021). Concerns about the accuracy and usefulness of bot detection have also been raised by security and integrity experts at major platforms (Roth and Pickles, 2020), as well as by independent researchers and the media. [^4] https://www.nytimes.com/2020/06/16/science/social-media-bots-kazemi.html [^5] https://mediawell.ssrc.org/expert-reflections/on-digital-disinformation-and-democratic-myths/ [^6] https://www.lawfareblog.com/random-toxicity-whats-going-benjaminwittess-mentions The growing tension between the continuing challenges of bot detection and the increasing technological capabilities of online manipulators, including unprecedented affordances such as those offered by generative AI (Yang and Menczer, 2023; Ferrara, 2023), even led some researchers to question the long-term viability of the bot detection task (Boneh et al., 2019; Grimme et al., 2022). For the same reasons, some have suggested redirecting part of the ongoing scientific effort away from bot detection and to alternative and more promising tasks, such as detecting information operations and coordinated malicious behavior (Cresci, 2020; Roth and Pickles, 2020; Pacheco et al., 2021). This sampling of the existing literature on the challenges and limitations of bot detection paints a very different picture than some recent, optimistic works. This discrepancy introduces a common misconception about social bots research:
@@
For many years, we have witnessed a game of whack-a-mole between bot developers, with their increasingly sophisticated accounts (Yang and Menczer, 2023; Yang et al., 2024), and bot hunters, equipped with a variety of different detectors (Yang et al., 2019; Sayyadiharikandeh et al., 2020; Yang et al., 2020; Cresci et al., 2021). Looking back at how this arms race has unfolded, we can conclude that none of the technological advances we have experimented with have significantly mitigated the challenges posed by malicious social bots in the long term. It is reasonable to assume that future advances will suffer a similar fate. Social bots are adversarial, fast-moving targets, characterized by a fast adoption of cutting-edge technology. Bot detection is therefore an intrinsically challenging task, made even more daunting by the lack of accurate information on the targets of the analysis, the limited collaboration from online platforms, and the rapidly evolving nature of online harms.
@@
An unbiased analysis of the existing literature paints a much more dubious picture of the role of social bots. For example, while some studies have concluded that bots play a prominent role in the spread of problematic content (Shao et al., 2018; Stella et al., 2018), others have reached the opposite conclusion (Vosoughi et al., 2018; González-Bailón and De Domenico, 2021; Seckin et al., 2024). The existing literature has focused almost exclusively on detecting bots and characterizing their behavior, leaving the fundamental task of measuring the impact of bot malfeasance largely unexplored (Cresci, 2020). For these reasons, we currently lack scientific consensus and conclusive evidence about the role of social bots and their effectiveness in influencing online users. What we do know, however, is that bots are only one of many agents involved in the spread of misinformation and disinformation (Starbird, 2019; Roth and Pickles, 2020; Pierri et al., 2023). Examples of some of the other agents include state-sponsored trolls, users who collude and coordinate for malicious purposes, superspreaders, and even willing but unwitting individuals (Starbird, 2019; DeVerna et al., 2022). Each of these agents represents a potential threat to safe and trusted online platforms, and a thriving area of research and experimentation. It is therefore critical to balance efforts in all of these directions, avoiding the pitfall of overusing some while overlooking others, based on unsupported decisions.
@@
We see at least two strong arguments against this thesis. First, despite the limitations of bot detectors, there are several glaring examples of bot studies that were able to bring to light demonstrably harmful campaigns. For example, detectors developed as part of some scientific efforts were later deployed on online platforms and used to remove large numbers of malicious accounts (Yang et al., 2014). Similarly, the results of some bot activity studies led platforms to remove accounts identified as malicious bots (Ferrara, 2022; Yang and Menczer, 2023). In several other cases, scientific findings about bot activity were later found to be largely consistent with independent platform removals of malicious accounts (Nizzoli et al., 2020; Tardelli et al., 2022), confirming the accuracy of the findings. These cases are just some of the success stories of social bot research. If researchers had not developed bot detection techniques, we would not have been able to identify pockets of anomalous accounts engaged in sophisticated manipulations. Therefore, even if no universal bot detector exists yet, and despite the many caveats to consider in bot detection, being able to detect some malicious bots puts us in a more advantageous position than being able to detect none. Second, the benefits of social bot research extend beyond the detection of malicious bots. For example, research and experimentation on social bots led to the development of neutral bots used to assess the level of political polarization and bias on a platform (Chen et al., 2021); “news bots” used for journalistic purposes to curate, aggregate, and distribute content gathered from multiple sources (Lokot and Diakopoulos, 2016); and even bots used for content moderation (Bilewicz et al., 2021; Askari et al., 2024). In addition, social bot research has contributed to the early development of other neighboring fields. The early research on social bots, which dates back to 2010 (Ratkiewicz et al., 2011; Cresci, 2020), provided an important foundation, allowing the field to draw upon several years of accrued experience with related phenomena, thereby offering crucial insights when widespread concerns regarding misinformation, state-backed trolls, and coordinated inauthentic behavior surged in 2016 and subsequent years. In other words, early results on detecting and characterizing social bots informed strategies for detecting and mitigating other related forms of online manipulation. In light of these considerations, research on social bots—imperfect as it is—seems far from useless. In a sense, research on social bots is no exception to a core framework for the advancement of science: The accumulation of knowledge as a result of research efforts, other than solving only the specific problems the research is addressed to, also fertilizes the scientific ecosystem in many ways (e.g., forming generations of scientists) and provides tools, data, methodology, and insights for scientific advancements in close (and sometimes not so close) research fields.
Challenges in the post-API era

### v3

@@ 2 Methodological issues
Unfortunately, some works in literature fail to follow these principles, resulting in inconsistent labeling and misleading results. Without disclosing any annotation rubric, Gallwitz and Kreil (2022) change criteria for labeling bots multiple times, thus inflating the false-positive or false-negative rates of bot detection methods to fit their narrative. In one case, criticizing Yang et al. (2019), they label accounts that automatically post news headlines—such as those associated with major newswire agencies—as human. Their criticism of Keller and Klinger (2019) goes further. First, Gallwitz and Kreil considered a different set of accounts than those analyzed by Keller and Klinger. Second, they manually inspected only 0.04% of this different set. Third, they ignored suspended accounts, and fourth, they labeled as human accounts those that cross-posted tweets through software apps. The aforementioned four arbitrary choices imply that they did not reproduce Keller and Klinger’s analysis. As such, their conclusion that «not a single one of the accounts in [Keller and Klinger’s] list was a “bot” in any meaningful sense of the word» is simply false. Similarly biased annotations are used by Gallwitz and Kreil (2022) to criticize Dunn et al. (2020), as suspended accounts are again assumed to be human to boost the alleged false-positive rate. Then, accounts that are clearly automated are still presented as classification errors by labeling them as “(unmalicious) bot” and accounts “spreading vaccine-critical information.” The use of obsolete or biased datasets with unclear, inconsistent, or shifting definitions are among the malpractices that affect this field.
Table 1: Summary of the main challenges and issues that may occur at each step of the social bots research process. For each issue, we propose a set of guidelines and practical recommendations. Since not all challenges apply to every study, it is up to the authors of future works to review this table to identify relevant issues specific to their study and follow the corresponding recommendations. The table could also serve as a valuable resource for reviewers and evaluators to assess the rigor of future works in terms of how new works implement the best practices and recommendations listed.
step issues recommendations
[…]
Practical recommendations
The above sections highlighted a number of challenges and widespread methodological issues that currently affect the field of social bots research. Table LABEL:tab:checklist summarizes the main methodological issues and proposes a set of guidelines and practical recommendations to mitigate said issues so as to produce more reliable and robust research. The main issues and recommendations are also discussed in the following.
One significant challenge is the use of contrasting definitions of social bots, which leads to incomparable results and confusion in research. To mitigate this, future studies should openly and clearly define the specific characteristics of the social bots they are targeting. Moreover, future research should adopt definitions consistently, without changing bot criteria and definitions case-by-case. Furthermore, favoring definitions based on objective characteristics—such as the use of automation—over subjective ones, will boost clarity and consistency. When discussing results, researchers should also compare their definitions with existing ones to highlight similarities and differences, also exploring how different definitions could have impacted the results. A related pressing issue is the use of inconsistent or opaque bot labeling procedures, as these hinder the clarity, comparability, and reproducibility of results. Researchers should document and publish their labeling procedures, including the criteria and tools used to carry out the annotation. Sharing labeled datasets with comprehensive metadata will further enhance transparency, reproducibility, and collaborative efforts.
A further set of recommendations involve the evaluation procedures of social bot detectors. Newly trained bot detectors are often tested on well-known and old bots, which does not accurately reflect their performance in real-world scenarios. Researchers should prioritize using the most up-to-date datasets for their evaluations, as these are likely to better reflect the evolving nature of social bots. Performance assessments should include tests on bots with different and unknown characteristics, to better gauge the detectors’ real-world applicability and to ensure that detectors are robust and effective in dynamic online environments. In addition, attention to validation should not only be the focus of the training and testing phases of a new bot detector, but should also be devoted when using a pre-trained detector developed by others, even if it is well-known and established. A subset of automatically assigned labels and scores should always be manually validated, rather than accepted at face value. This ensures their accuracy and reliability in the given context, which might be different than that on which the detector was developed. As with the aforementioned manual labeling procedures, also this manual validation process should be clearly and openly documented, detailing the criteria and methods used for manual verification of the assigned labels.
Another issue in the evaluation of bot detection performance is the possibility to cherry-pick competitors in such a way to purportedly demonstrate the superior performance of a newly proposed detector. To mitigate this issue, future research should benchmark bot detectors against a comprehensive set of existing detectors, including both strong and weak performers. Moreover, the criteria for the selection of the comparisons should be stated clearly and openly. Sharing evaluation scripts and datasets used in benchmarking will further promote transparency and reproducibility in performance comparisons. In addition, sensitivity analyses should also become a standard practice to address the possible sensitivity of detectors to thresholds and other model parameters, which are often overlooked when reporting bot detection results. These analyses should be reported alongside the main results to provide a clearer picture of the robustness and reliability of the detectors, aiding in the identification of optimal settings and limitations.
Lastly, data availability and documentation are crucial for the progress of bot detection research. The lack of shared and well-documented datasets impedes reproducibility and comparison, and slows down the development of new detectors. Researchers should make their datasets publicly available whenever possible, ensuring they are well-documented with comprehensive metadata, including how the data was collected and labelled. This recommendation does not only involve the datasets used to train or evaluate a detector, such as those labeled manually, but also the datasets on which pre-trained detectors are applied. For the latter, it is crucial to share the data together with the labels or scores automatically assigned by the detectors. If data sharing is constrained by privacy or other issues, researchers should at least provide detailed descriptions of their datasets and the procedures for obtaining them. This practice will support reproducibility and facilitate more rigorous comparisons and advancements in bot detection research.
@@ 3 Conceptual issues
The assumption about the possibility to build comprehensive bot datasets is equally flawed. Creating such datasets would require a uniform and random sample of an adequate number of accounts from the complete distribution of existing bots. This would allow obtaining an accurate representation of the bot landscape that captures the full complexity of the problem space. While in some domains of application for machine learning this is feasible, [^2] For example, in certain speech recognition tasks all instances (e.g., languages) to be classified are known in advance. Given a sufficiently large and diverse set of user inputs, unbiased and comprehensive datasets are possible (Ardila et al., 2020). we argue that performing a uniform and random sample of the whole population of bots is extremely problematic—if not outright impossible. In addition to the many general challenges brought by social data (Olteanu et al., 2019; Pasquetto et al., 2020), the task of detecting social bots also partly belongs to the fields of information security and open-source intelligence, which are intrinsically characterized by the presence of adversaries who are strongly motivated to remain hidden (Cresci et al., 2021). Indeed, there has always been limited knowledge of the real extent of the bot problem, which makes it hard to track bots on a platform, quantify their numbers, and assess their impact (Varol, 2023; Mendoza et al., 2020; Tan et al., 2023). We only have a partial understanding and visibility of the botnets that operate on our online platforms. [^3] The court dispute between Elon Musk and former Twitter CEO Parag Agrawal on the real number of bots on Twitter is a paramount example of the intrinsic difficulties and ambiguity of tracking social bots. Figure 1 provides a conceptual view of the landscape of social bots and human-operated accounts: firstly according to a naïve interpretation (Figure 1(a)) and then in a more realistic representation (Figure 1(b)). Figure 1(b) highlights the knowledge gaps related to the presence of bots in online platforms. These include known unknowns: the exact number of bots belonging to known botnets. But they also include unknown unknowns: information about the hidden botnets operating on a platform. How can we create unbiased and comprehensive datasets when we largely ignore the types and numbers of bots populating online platforms? This fundamental question cannot currently be answered.
The previous discussion highlights the challenges in obtaining a reliable ground-truth for bot detection, which is largely due to the impossibility of acquiring accurate and dependable bot labels. Researchers are often compelled to use proxies, such as platform account removals or manual annotations based on profile characteristics, which however reduce the reliability and validity of the ground-truth. Furthermore, the lack of comprehensive knowledge about the types, sizes, and characteristics of bots operating on platforms creates an imbalance where data about certain bot types may be more accessible than others. Data availability can facilitate research on specific sub-types of bots, potentially leading to more favorable results, albeit confined to a limited subset of bots. However, this could also introduce bias in the development of bot detectors, making them particularly effective against the overrepresented bot types while being ineffective against the underrepresented ones.
@@
Some recent work has interpreted the results of social bot detection studies as a success story for machine learning, concluding that the positive results seem to suggest that the bot detection task is now effectively solved (Hays et al., 2023). Proponents of this perspective highlight the perceived accuracy and reliability of current detection methods, and advocate their suitability for integration into various downstream applications, highlighting the progress made in achieving practical and deployable solutions in the field.
@@
We disagree with this interpretation of the state of social bot research. On the contrary, we believe that unbiased and responsible research in this area should acknowledge the body of literature warning researchers and practitioners about the limitations of bot detection. We are well aware of this body of work, having actively contributed to its development for many years. Below we report on some notable examples. The first accounts of the difficulties supervised detectors face in detecting sophisticated bots date back to 2013 (Yang et al., 2013) and intensified in subsequent years (Cresci et al., 2017; Grimme et al., 2018; Cresci et al., 2019b). The limitations of benchmark datasets for bot detection were discussed by Feng et al. (2021) and the data-related limitations to the generalizability and replicability in social bots research were touched upon by Assenmacher et al. (2022). Several works measured and discussed the generalization deficiencies of bot detectors, with particular emphasis on supervised ones (Echeverría et al., 2018; Yang et al., 2019; Cresci, 2020; Sayyadiharikandeh et al., 2020; Yang et al., 2020; Feng et al., 2021). Others warned about the lack of consistency in bot classifications obtained at different points in time (Rauchfleisch and Kaiser, 2020) and with different detectors (Svenaeus, 2020; Martini et al., 2021). Concerns about the accuracy and usefulness of bot detection have also been raised by security and integrity experts at major platforms (Roth and Pickles, 2020), as well as by independent researchers and the media. [^4] https://www.nytimes.com/2020/06/16/science/social-media-bots-kazemi.html [^5] https://mediawell.ssrc.org/expert-reflections/on-digital-disinformation-and-democratic-myths/ [^6] https://www.lawfareblog.com/random-toxicity-whats-going-benjaminwittess-mentions The growing tension between the continuing challenges of bot detection and the increasing technological capabilities of online manipulators, including unprecedented affordances such as those offered by generative AI (Yang and Menczer, 2023; Ferrara, 2023), even led some researchers to question the long-term viability of the bot detection task (Boneh et al., 2019; Grimme et al., 2022). For the same reasons, some have suggested redirecting part of the ongoing scientific effort away from bot detection and to alternative and more promising tasks, such as detecting information operations and coordinated malicious behavior (Cresci, 2020; Roth and Pickles, 2020). This sampling of the existing literature on the challenges and limitations of bot detection paints a very different picture than some recent, optimistic works. This discrepancy introduces a common misconception about social bots research:
@@
For many years, we have witnessed a game of whack-a-mole between bot developers, with their increasingly sophisticated accounts (Yang and Menczer, 2023; Yang et al., 2024), and bot hunters, equipped with a variety of different detectors (Yang et al., 2019; Sayyadiharikandeh et al., 2020; Yang et al., 2020). Looking back at how this arms race has unfolded (Cresci et al., 2021), we can conclude that none of the technological advances we have experimented with have significantly mitigated the challenges posed by malicious social bots in the long term. It is reasonable to assume that future advances will suffer a similar fate. Social bots are adversarial, fast-moving targets, characterized by a fast adoption of cutting-edge technology. Bot detection is therefore an intrinsically challenging task, made even more daunting by the lack of accurate information on the targets of the analysis, the limited collaboration from online platforms, and the rapidly evolving nature of online harms.
@@
An unbiased analysis of the existing literature paints a much more dubious picture of the role of social bots. For example, while some studies have concluded that bots play a prominent role in the spread of problematic content (Shao et al., 2018; Stella et al., 2018), others have reached the opposite conclusion (Vosoughi et al., 2018; González-Bailón and De Domenico, 2021; Seckin et al., 2024). The existing literature has focused almost exclusively on detecting bots and characterizing their behavior, leaving the fundamental task of measuring the impact of bot malfeasance largely unexplored (Cresci, 2020). For these reasons, we currently lack scientific consensus and conclusive evidence about the role of social bots and their effectiveness in influencing online users. What we do know, however, is that bots are only one of many agents involved in the spread of misinformation and disinformation (Starbird, 2019; Roth and Pickles, 2020). Examples of some of the other agents include state-sponsored trolls, users who collude and coordinate for malicious purposes, superspreaders, and even willing but unwitting individuals (Starbird, 2019; DeVerna et al., 2022). Each of these agents represents a potential threat to safe and trusted online platforms, and a thriving area of research and experimentation. It is therefore critical to balance efforts in all of these directions, avoiding the pitfall of overusing some while overlooking others, based on unsupported decisions.
@@
reference main result validation or corroboration source
Yang et al. (2014) studied social bots and developed effective detection techniques platform collaborations in real-world scenarios platform deployments in production environments
Subrahmanian et al. (2016) hosted and presented results of the DARPA Twitter bot challenge findings incorporated into defense and cybersecurity strategies U.S. Department of Defense reports and cybersecurity policies
Ferrara et al. (2016) brought to light the issue of social bots and their capabilities in manipulation confirmed from various data leaks showing coordinated bot operations federal investigation reports, media disclosures
Cresci et al. (2017) demonstrated that social bots evolve, becoming more sophisticated with time supported by subsequent studies and data showing similar patterns academic follow-up studies and industry research
Varol et al. (2017) estimated social bots prevalence reports and admissions by social media platforms public statements and transparency reports from social media
Lazer et al. (2018) highlighted the impact of bots on misinformation spread and public opinion alignment with findings from investigations into electoral interference government investigation documents and hearings
Badawy et al. (2019) analyzed use of bots in Russian interference with the 2016 U.S. election corroborated by U.S. intelligence community findings U.S. Senate and House intelligence committee reports
Cresci (2020) framed social bot detection as an adversarial machine learning task demonstrated by the improvements obtained via adversarial approaches academic follow-up studies and industry research
Ferrara (2022) provided a taxonomy of Twitter bots, methods, and tools to detect them results supported by platforms and public organizations reports and investigations by government agencies
Table 2: Examples of key findings on social bots that have withstood the test of time. Each result is accompanied by details on its validation or corroboration, along with the supporting sources. While not exhaustive, this list underscores the impact that rigorous social bots research can achieve.
@@
We see at least three strong arguments against this thesis. First, despite the limitations of bot detectors, there are several glaring examples of bot studies that were able to bring to light demonstrably harmful campaigns. For example, detectors developed as part of some scientific efforts were later deployed on online platforms and used to remove large numbers of malicious accounts (Yang et al., 2014). Similarly, the results of some bot activity studies led platforms to remove accounts identified as malicious bots (Ferrara, 2022; Yang and Menczer, 2023). In several other cases, scientific findings about bot activity were later found to be largely consistent with independent platform removals of malicious accounts (Nizzoli et al., 2020; Tardelli et al., 2022), confirming the accuracy of the scientific findings. These cases are just some of the success stories of social bot research. Without the development of bot detection techniques, identifying clusters of anomalous accounts engaged in sophisticated manipulations would not have been possible. Therefore, even if no universal bot detector exists yet, and despite the many caveats to consider in bot detection, being able to detect some malicious bots puts us in a more advantageous position than being able to detect none.
Second, the methodological rigor and practical usefulness of some social bots studies are evidenced by corrective actions taken by government agencies based on academic research. For instance, in November 2014, the U.S. Securities and Exchange Commission (SEC) issued an alert to raise awareness about stock market manipulations exacerbated by social bots, [^7] Updated Investor Alert: Social Media and Investing – Avoiding Fraud (Nov. 12, 2014) https://www.sec.gov/resources-for-investors/investor-alerts-bulletins/ia_socialmediafraud following the findings of earlier studies (Hwang et al., 2012). Subsequent research confirmed these suspicions in multiple occasions (Cresci et al., 2019a; Nizzoli et al., 2020), contributing to calls for greater government oversight of social bots (Gorwa and Guilbeault, 2020a; Yan et al., 2023) and discussions on regulating bot freedom of expression (Lamo and Calo, 2019). Both the 2018 and 2022 EU Code of Practice on Disinformation, as well as investigations by the U.S. Congress and the UK Parliament’s Digital, Culture, Media and Sport Committee have also been informed by research on social bots. Furthermore, the concern over bot infiltration was highlighted when Elon Musk initially decided to abandon his purchase of Twitter, citing misleading information about the platform’s financial health and bot prevalence (Varol, 2023)—a claim also supported by Peiter Zatko, Twitter’s former head of security, who exposed serious deficiencies in the company’s security. [^8] What the Twitter Whistleblower Disclosure Means for Elon Musk: https://time.com/6207992/twitter-bots-elon-musk/ Table 2 provides some examples of contributions in the study and mitigation of social bots that have withstood the test of time. These examples underscore the enduring impact and significance of rigorous social bots research in shaping policy and industry practices.
Third, the benefits of social bot research extend beyond the detection of malicious bots. For example, research and experimentation on social bots led to the development of neutral bots used to assess the level of political polarization and bias on a platform (Chen et al., 2021); “news bots” used for journalistic purposes to curate, aggregate, and distribute content gathered from multiple sources (Lokot and Diakopoulos, 2016); and even bots used for content moderation (Bilewicz et al., 2021; Askari et al., 2024). In addition, social bot research has contributed to the early development of other neighboring fields. The early research on social bots, which dates back to 2010 (Cresci, 2020), provided an important foundation, allowing the field to draw upon several years of accrued experience with related phenomena, thereby offering crucial insights when widespread concerns regarding misinformation, state-backed trolls, and coordinated inauthentic behavior surged in 2016 and subsequent years. In other words, early results on detecting and characterizing social bots informed strategies for detecting and mitigating other related forms of online manipulation. In light of these considerations, research on social bots—imperfect as it is—seems far from useless. In a sense, research on social bots is no exception to a core framework for the advancement of science: The accumulation of knowledge as a result of research efforts, other than solving only the specific problems the research is addressed to, also fertilizes the scientific ecosystem in many ways (e.g., forming generations of scientists) and provides tools, data, methodology, and insights for scientific advancements in close (and sometimes not so close) research fields.
4 Challenges in the post-API era